OpenStudy (anonymous):

Solve the system 2x - 3y = 0 and x + 4y = -11 by graphing each line.

OpenStudy (anonymous):

well both equations are lines so you can easily graph them by plotting some points. but the best way probably is to see how that equation intersect with the x axis, y axis, and each other lines *x-intersection (y=0) for example for 2x-3y=0 plotting y=0 into the equation will give 2x-0=0, which will then prove that x=0 *y-intersection(x=0) the same way with the above *intersection with each other equation: basically set them equal to each other. from the first equation we get 2x=3y -> x=3/2y plugging this to the second equation will give you the points of intersection, which is supposed to be the solution for your answer